OpenStudy (anonymous):

*MEDAL* Show your work to calculate the discriminant of the quadratic equation, and then state the number and type of solutions you expect the equation to have, based on the value of your discriminant. 0=x^2+6x-16

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

okay I worked it out on my own and got this. (6)^2-4(1)(-16) 36-4(-16) -4x-16=64 4+64=68 I think I messed up somewhere but I'm not sure where.

OpenStudy (anonymous):

where did the x come from?

OpenStudy (anonymous):

\[(6)^2-4(1)(-16)\] is right

OpenStudy (anonymous):

\[36-4(-16)\] is right too

OpenStudy (anonymous):

some gibberish after than where in the heck the the x come from?

OpenStudy (anonymous):

what is \(-4\times (-16)\)?

OpenStudy (anonymous):

I meant -4(-16)

OpenStudy (anonymous):

right and what is that number?

OpenStudy (anonymous):

64

OpenStudy (anonymous):

ok and what is \(36+64\) ?

OpenStudy (anonymous):

100. The video my teacher posted got me confused when she added the last numbers because I thought she just added 4 to 64 and it only left me with 68

OpenStudy (anonymous):

no it is 100 for sure